a research center claims that 25% A ro con c claims that 25 could afford it l a pe o 1200 aduits that country. 28% say tha they would ra d c cal f they could a
Answer #1

a) Option - C) 25% of adults in the country would travel into space on a commercial flight if they could afford it.

Option - F) H0: P = 0.25

                  H1: P \neq 0.25

b) The test statistic z = (\widehat p - P)/sqrt(P(1 - P)/n)

                                 = (0.28 - 0.25)/sqrt(0.25 * (1 - 0.25)/1200)

                                 = 2.4

P-value = 2 * P(Z > 2.4)

            = 2 * (1 - P(Z < 2.4))

            = 2 * (1 - 0.9918)

            = 0.0164 = 0.016

c) Since the P-value is greater than the significance level(0.016 > 0.01), so we should not reject the null hypothesis.

Do not reject the null hypothesis. There is enough evidence to support the research center's claim.